# Open, Monetizable, and Loyal AI

### v1

@@ 2.1.2 Construction and Security
    Hardness of generating the permission strings $\sigma(h(x))$ without the secret key.
Satisfying both conditions ensures model loyalty through OML formatting. More rigorously, assume an adversary has some set of inputs $\{x_{i}\}_{i=1}^{n}$, a set of permission string instances $\{(h(x_{i}),\sigma(h(x_{i}))\}_{i=1}^{n}$, and hence corresponding model outputs, at a cost proportional to the number of samples $n$, legitimately acquired by using the OML file. Potential attacks to get continual and permissionless access to the OML-formatted model include several strategies. An attacker can analyze the dynamics of the OML-formatted model, for example, using techniques from neural network surgery [27, 28] and engineer how to re-wire the model to bypass the verification of the permission string. An attacker can also use the labelled data, $\{(h(x_{i}),\sigma(h(x_{i}))\}_{i=1}^{n}$, to recover the mapping $\sigma(\cdot)$. The goal of an ideal OML primitive is to make it as costly as possible for the attacker to launch such attacks, using a combination of cryptographic, statistical, and machine learning tools.
2.2 Canonical OML Constructions
In this section, we will discuss different possible approaches to OML formatting, their security and performance implications. We will discuss them in order of ascending strength of security guarantees, followed by a melange construction that can provide the most flexibility in defining the most desirable OML format for various model owners.
@@ 2.2 Canonical OML Constructions
    Obfuscation [Software security]. Software obfuscation is a set of methods that reformat a program $P$ into a functionally equivalent yet hard-to-understand program $P^{\prime}$, where $P(x)=P^{\prime}(x)$ for all inputs $x$. Through obfuscation techniques, including optimized compilation, an OML formatted model is much harder for adversaries to analyze, understand, and therefore modify, providing protection against model stealing.
@@
    Fingerprinting [Optimistic security]. We describe this OML method as optimistic OML, a novel monetizable mechanism based on data poisoning techniques. More specifically, we plant several (pre-defined input, expected output) pairs that act as backdoors on the model such that its ownership can be verified after it is distributed to the users. In a sense, optimistic OML uses the same idea as optimistic rollups from blockchains where any deployed model is assumed to be non-stolen unless challenged. The validity of the challenge is then verified, and the malicious user can be punished appropriately.
@@
    Trusted execution environments (TEEs) [Hardware security]. A TEE is a hardware-enabled enclave that can run arbitrary code on an untrusted machine without exposing any code to the machine host. An AI model would be downloaded by a user in encrypted format and only be decrypted within the TEE. The security of a TEE is thus reliant on the vendor of the corresponding hardware and possible implementation-specific jailbreaks, with Intel SGX/TDX, AMD SEV, and Arm TrustZone being the corresponding implementations by the largest vendors. While there is overhead from encryption, decryption, and secure channel communication between the TEE and the untrusted host, actually running programs inside the TEE incurs no extra performance cost compared to running them on the untrusted host as usual. TEEs can also be scaled up to the machine’s near maximum available resources, allowing for creation of very large enclaves to hold giant AI models. The main limitation of TEEs is that only CPU TEEs are commercially available at the moment, imposing limitations on what kinds of AI workloads can be done efficiently on local hardware-enabled enclaves.
@@
    Cryptography [Provable security]. The strongest security is based on impossibility results backed by provable cryptographic hardness, and can be achieved by state-of-the-art cryptographic primitives such as Fully Homomorphic Encryption (FHE). It can be theoretically shown that no adversary can break an FHE-based OML file unless some unlikely fundamental assumptions (e.g. hardness of well-known problems like lattices) are compromised. This level of security often comes with significant performance and processing overhead, and is suitable for models that have high monetary value and are small or do not expect high throughput or frequent usage.
@@ 5. $\star$
    Melange – mixed OML. The aforementioned methods can be combined and methodically applied to the entire model or separate parts of the model. In this way, OML can adapt to the needs of different model owners, enabling flexible security guarantees. With this approach, model owners are given full control of how their model is separated, how different security methods are used, and what combination of these methods defines their own preferred version of OML.
In the following sections, we will go into more detail on these OML approaches and discuss what the OML formatting phase and the verification and usage phase look like for each approach. For each, we will identify which characteristics of open models (transparency, locality, mutability and privacy) are sacrificed and to what extent.
2.2.1 Obfuscation
@@
By combining all these techniques, we can come up with a clear construction for OML (Figure 2.2).
Figure 2.2: OML formatting process of AI models via obfuscation.
OML formatting. Recall, from Section 2.1.2 that a naive OML file can be constructed simply by prepending the permission string verification function ${\rm Verify}_{\rm pk}$ to the plain-text model $M$, with the model only returning the correct result if the verification passes. This implies that an attacker can easily find and remove the verification function in the code, recovering the use of the model without the need for permission. To safeguard this OML construction, software obfuscation techniques can be applied such that the two components (${\rm Verify}_{\rm pk}$ and $M$) are intermingled with one another, represented as non-comprehensible code with complicated control flow. As a result, it is difficult to pinpoint the exact location of ${\rm Verify}_{\rm pk}$ in the obfuscated OML file, making it hard for an attacker to remove verification and recover the original model $M$.
@@
2.2.2 Fingerprinting
Optimistic OML prioritizes efficiency while ensuring a weaker notion of next-day security, i.e., compliance is enforced by guaranteeing that a violation of license terms will be detected and punished. Inspired by optimistic security [26], optimistic OML relies on compliance with the license, and compensating transactions are used to ensure that the model owners’ rights are protected, in case of a violation. Crucial in this process are techniques for authenticating the ownership of a model. For example, Llama models [23] are released under a unique license that a licensee with more than 700 million monthly active users is “not authorized to exercise any of the rights under this Agreement unless or until Meta otherwise expressly grants you such rights”. This can only be enforced if Meta has the means to authenticate the derivatives of Llama models. We propose planting a backdoor on the model such that it memorizes carefully chosen fingerprint pairs of the form (key, target response). If successful, such fingerprints can be checked after deployment to claim ownership. An optimistic OML technique should satisfy the following criteria:

### v3

@@ 2.1 The OML Protocol and Design Space
    Permission Scheme ($\sigma,h$): The functions $h:\mathcal{X}\to\mathcal{H}$ and $\sigma:\mathcal{H}\times\mathcal{K}_{own}\to\mathcal{P}$ that define how permission tokens are generated and bound to inputs. The security of this scheme (e.g., unforgeability of $p_{x}$) is paramount.
  3. 3.
    OMLization Process ($OMLize$): A transformation $M_{oml}\leftarrow OMLize(M,\text{params}_{\sigma,h})$. This process embeds the logic within $M_{oml}$ to verify $p_{x}$ and conditions its correct execution on this verification. $\text{params}_{\sigma,h}$ might include public parameters of the permission scheme (e.g., a public verification key $vk_{own}$ corresponding to $k_{own}$).
2.2 Core Properties of the OML Format
Let $M:\mathcal{X}\to\mathcal{Y}$ be a plain-text AI model that we want to protect, where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the output space. The OML cryptographic primitive transforms $M$ into an OML-formatted model, $M_{oml}$, designed to enable its distribution while protecting the intellectual property (IP) rights of the model owner, denoted $\mathcal{O}$. This balance ensures that AI models can be monetized effectively and remain loyal to $\mathcal{O}$.
Property 2.1 (Open-access (O)).
The Open-access (O) property signifies that $M_{oml}$ is openly accessible to everyone which provides the following specific operational benefits:
  • (a)
    Local Hosting and Execution: $M_{oml}$ is available for download (e.g., via a distribution channel $\mathcal{D}$), allowing users to host and execute it on their local infrastructure. This facilitates workflow optimization, can enhance data privacy for the user by keeping inputs local, and ensures consistent service quality independent of $\mathcal{O}$’s serving infrastructure.
  • (b)
    Local Enhancements: Users can, upon approval from the model owner $\mathcal{O}$, apply standard model improvement techniques locally to their copy of $M_{oml}$ (e.g., fine-tuning with local data, retrieval-augmented generation).
  • (c)
    Integrity and Post-Publication Immutability: The open accessibility of an $M_{oml}$ artifact provides strong assurances of its integrity. Specifically, once $M_{oml}$ is published and distributed, the model owner $\mathcal{O}$ cannot unilaterally alter that particular instance. This immutability by the owner ensures that users are interacting with the model as it was originally released, fostering trust in the specific version they possess.
Property 2.2 (Monetizable (M)).
The Monetizable (M) property ensures that an OML-formatted model $M_{oml}$ enables its owner $\mathcal{O}$ to derive economic value for each usage through a fine-grained, per-instance authorization protocol. The core mechanisms are detailed as follows:
  • (a)
    Granular Authorization: Access to the utility of $M_{oml}$ for a specific input $x\in\mathcal{X}$ requires a valid permission token $p_{x}$. This token is issued or verifiable by $\mathcal{O}$ and is essential for implementing flexible, pay-per-use billing models (e.g., per-input or per-token charges). This granular approach contrasts sharply with typical open-weight distributions where monetization strategies are often restricted to coarse-grained, model-level access (e.g., a one-time purchase for full model download or a general license for unlimited use).
  • (b)
    Conditional Value Proposition: Let $U(M,x)$ denote the utility derived from the output $M(x)$ of the original plain-text model. The OML model $M_{oml}$ ensures that a user $\mathcal{U}$ can only achieve utility comparable to $U(M,x)$ by adhering to $\mathcal{O}$’s terms for that specific usage instance. This adherence centrally involves acquiring and utilizing the valid permission token $p_{x}$, as unauthorized attempts to use $M_{oml}$ should not yield the desired utility.
Property 2.3 (Loyal (L)).
Loyalty ensures that an OML-formatted model $M_{oml}$ strictly adheres to its owner $\mathcal{O}$’s operational terms, primarily through the pre-hoc enforcement of authorization. This pre-hoc control is a crucial mechanism for addressing AI safety and alignment concerns by enabling owner-defined governance over the model’s use.
  • (a)
    Proof-of-Ownership: The owner $\mathcal{O}$ possesses a secret key $k_{own}\in\mathcal{K}_{own}$, which is fundamental to the generation or verification of permission tokens $p_{x}$ required for authorized use. The secret key can be either AI-native or cryptographic depending on the specific scheme.
  • (b)
    Pre-hoc Authorization Enforcement: $M_{oml}$ is engineered such that it will not produce a correct or high-utility output for an input $x$ unless a valid permission token $p_{x}$ (derived using $k_{own}$) is provided as part of its input. This authorization must occur before $M_{oml}$ completes its useful computation. Such enforcement is critical for high-value models and for scenarios requiring diligent control over AI safety and alignment, as it proactively prevents unauthorized or non-compliant content generation.
  • (c)
    Controlled Openness: While $M_{oml}$ itself might be openly accessible (as per Property 2.1), its correct and intended functioning is inextricably tied to $\mathcal{O}$’s authorization mechanisms. This owner-centric control can be designed to be auditable if the authorization protocol itself is transparent (e.g., through rules managed by a smart contract).
These three properties collectively define the desired behavior of an OML model. Crucially, the OML format must guarantee that unauthorized users cannot circumvent these IP rights and access controls to gain unpaid or unrestricted access to $M$’s capabilities through $M_{oml}$ without incurring prohibitive computational expense.
2.3 Security Guarantees and Adversarial Model
The primary security objective of the OML primitive is to enforce Properties 2.2 (Monetizability) and 2.3 (Loyalty), even when an adversary $\mathcal{A}$ has white-box access to $M_{oml}$.
Adversary Definition and Capabilities.
An adversary $\mathcal{A}$ is a computationally bounded algorithm whose goal is to circumvent the OML protections. $\mathcal{A}$ is assumed to possess:
  • Full white-box access to $M_{oml}$ (i.e., its architecture and parameters).
  • Access to an oracle $\mathcal{P}_{\mathcal{O}}(\cdot)$ that, for a chosen input $x_{i}\in\mathcal{X}$, returns the legitimate permission token $p_{x_{i}}=\sigma(h(x_{i}),k_{own})$. $\mathcal{A}$ can make up to $N$ queries to $\mathcal{P}_{\mathcal{O}}(\cdot)$, potentially at some cost per query. From these, $\mathcal{A}$ obtains a set of authorized input-token-output tuples $\mathcal{D}_{known}=\{(x_{i},p_{x_{i}},y_{i})\}_{i=1}^{N}$, where $y_{i}=M_{oml}(x_{i},p_{x_{i}})$.
Core Security Requirements (Hardness Properties).
We formally define the security of an OML scheme as follows. A secure OML scheme should satisfy both requirements 2.1 and 2.2.
Requirement 2.1 (Model Extraction Resistance).
It must be computationally infeasible for an adversary $\mathcal{A}$, given $M_{oml}$ and $\mathcal{D}_{known}$, to construct a standalone model $M^{\prime}$ that effectively replicates the functionality of the original model $M$ without requiring the OML permission mechanism. Consider the following experiment, $Expt_{\mathcal{A}}^{ME}$:
@@ 2.3 Security Guarantees and Adversarial Model
    $\mathcal{A}$ is given $M_{oml}$ and oracle access to $\mathcal{P}_{\mathcal{O}}(\cdot)$ for $N$ queries.
@@
    $\mathcal{A}$ outputs a model $M^{\prime}$.
@@
    $\mathcal{A}$ wins if, for a fresh input $x^{*}\sim\mathcal{D}_{\mathcal{X}}$ (drawn from the target distribution $\mathcal{D}_{\mathcal{X}}$, and $x^{*}\notin\{x_{i}\}_{i=1}^{N}$), $d(M^{\prime}(x^{*}),M(x^{*}))\leq\epsilon_{utility}$ with high probability.
The OML scheme is $(t,N,\epsilon_{ME})$-extraction resistant if for any adversary $\mathcal{A}$ running in time complexity at most $t$ and making $N$ queries, $Pr[Expt_{\mathcal{A}}^{ME}\text{ wins}]\leq\epsilon_{ME}$.
Requirement 2.2 (Permission Forgery Resistance).
It must be computationally infeasible for an adversary $\mathcal{A}$, given $M_{oml}$ and $\mathcal{D}_{known}$, to generate a valid permission token $p^{*}_{x^{*}}$ for a new input $x^{*}\notin\{x_{i}\}_{i=1}^{N}$ that allows $M_{oml}$ to produce a correct output for $x^{*}$. Consider the following experiment, $Expt_{\mathcal{A}}^{PF}$:
  1. 1.
    $\mathcal{A}$ is given $M_{oml}$ and oracle access to $\mathcal{P}_{\mathcal{O}}(\cdot)$ for $N$ queries.
  2. 2.
    A challenge input $x^{*}\sim\mathcal{D}_{\mathcal{X}}$ ($x^{*}\notin\{x_{i}\}_{i=1}^{N}$) is given to $\mathcal{A}$.
  3. 3.
    $\mathcal{A}$ outputs a permission token $p^{*}_{x^{*}}$.
@@
    $\mathcal{A}$ wins if $d(M_{oml}(x^{*},p^{*}_{x^{*}}),M(x^{*}))\leq\epsilon_{utility}$.
The OML scheme is $(t,N,\epsilon_{PF})$-forgery resistant if for any adversary $\mathcal{A}$ running in time at most $t$ and making $N$ queries, $Pr[Expt_{\mathcal{A}}^{PF}\text{ wins}]\leq\epsilon_{PF}$. This implies that $\mathcal{A}$ cannot recover $k_{own}$ or find a practical way to compute $\sigma(h(x^{*}),k_{own})$ without $k_{own}$.
Here, $\epsilon_{ME}$ and $\epsilon_{PF}$ should be negligible functions of the security parameter. Satisfying both requirements is essential for robustly ensuring loyalty and monetizability.
Threat Model: Potential Attack Strategies. Adversaries may employ several strategies to violate the above security requirements:
  1. 1.
    Verification Bypass or Removal: $\mathcal{A}$ analyzes $M_{oml}$ (e.g., using techniques from neural network surgery [85, 36] or program analysis) to identify and excise or neutralize the embedded permission verification logic. A successful attack would yield a model $M^{\prime}$ that behaves like $M$ but without requiring valid tokens, thus violating Requirement 2.1.
  2. 2.
    Verification Tampering: $\mathcal{A}$ attempts to manipulate the internal state or execution flow of $M_{oml}$ at runtime to force the verification step to succeed illicitly. This could lead to violating Requirement 2.2 for specific inputs if a generic bypass is not achieved.
  3. 3.
    Permission Forgery or Counterfeiting: $\mathcal{A}$ aims to produce valid $(x^{*},p^{*}_{x^{*}})$ pairs without legitimate access to $\mathcal{P}_{\mathcal{O}}(\cdot)$ for $x^{*}$. This includes:
    • (a)
      Directly crafting a $p^{*}_{x^{*}}$ that passes verification by exploiting weaknesses in $\sigma$ or its implementation within $M_{oml}$.
    • (b)
      Recovering $k_{own}$ or learning a functional equivalent of $\sigma(\cdot,k_{own})$ from $\mathcal{D}_{known}$.
    • (c)
      Training a surrogate model $M_{surrogate}(x)\approx p_{x}$ using $\mathcal{D}_{known}$ as training data.
Addressing Naive Constructions. A naive approach, such as $M_{oml}(x,p_{x}):=\text{If }Verify_{vk_{own}}(h(x),p_{x})\text{ then }M(x)\text{ else }\perp$, where $Verify$ is a standard digital signature verification algorithm and $p_{x}=Sign_{k_{own}}(h(x))$, is insecure. If $Verify$ and $M$ are loosely coupled components within $M_{oml}$, an adversary with white-box access can often easily identify and remove the $Verify$ call, effectively obtaining $M$. The $OMLize$ process must therefore deeply entangle the model’s computation with the verification logic.
The overarching goal of an OML primitive is to ensure that the computational cost for $\mathcal{A}$ to succeed in any of these attacks significantly outweighs the value derived, or that the success probability is negligibly small. This requires a sophisticated combination of cryptographic techniques, secure programming, and potentially obfuscation or machine learning-based defenses. By establishing these hardness properties, OML aims to create a fair and balanced ecosystem for AI models, protecting the rights of model owners without unduly restricting legitimate users.
3 Road to OML: Preliminary Solutions and Frameworks
This section outlines preliminary solutions and frameworks toward achieving robust Open-access, Monetizable, and Loyal (OML) model serving. We first explore theoretical OML constructions. We then discuss a practical instantiation, OML 1.0, based on AI-native model fingerprinting, alongside a proposed end-to-end framework for hosting and managing OML-formatted models. Detailed analyses for these topics are in Appendices A, B, and C.
3.1 OML in Theory - Canonical OML Construction Ideas
Various approaches can realize OML formatting, differing in their security guarantees and performance implications. These constructions, presented broadly by increasing security strength, include:
  1. 1.
    Obfuscation [Software security]: Leverage multi-layered software obfuscation techniques to protect models by making them hard to analyze and modify.
  2. 2.
    Trusted Execution Environments (TEEs) [Hardware security]: Leverage hardware-enforced secure enclaves to protect models during execution on remote untrusted machines.
  3. 3.
    Cryptography [Provable security]: Leverage primitives like Fully Homomorphic Encryption (FHE) for the strongest and provable security guarantees, often with performance trade-offs.
@@ 4. $\star$
    Melange – mixed OML: Combine the aforementioned methods flexibly to tailor security and performance to specific needs, applying different techniques to different model parts.
These construction-based OML candidates often come with efficiency overheads (especially for TEEs and advanced Cryptography), and their application can be strategic. They can serve as robust backup mechanisms for dispute resolution or challenge verification in hybrid systems that primarily rely on more efficient API-based black-box access, with the OML-protected model acting as the ultimate source of truth when users are unsatisfied with the provided output and challenges arise.
Detailed discussions of these approaches, including their respective formatting and verification phases, characteristic trade-offs, and a comprehensive comparative analysis, can be found in Appendix A.
3.2 OML in Practice: OML 1.0 with AI-Native Fingerprinting and Ecosystem Framework
As a first practical step towards OML, OML 1.0 utilizes AI-native cryptography, specifically model fingerprinting. This technique repurposes data poisoning methods to embed unique, secret (key, response) pairs into models. A key challenge in this domain is the design of fingerprints that are robust against removal or detection, even when an adversary has white-box access to the model and its entire inference process. The workflow of OML 1.0, optimistic OML, is shown in Figure 3.
Figure 3: Illustration of OML 1.0 Workflow
The OML 1.0 system operates on crypto-economic principles: model users (e.g., services hosting the model publicly) agree to usage terms and post collateral. Authorized provers can then query these hosted models with the secret keys. If a model user provides a fingerprinted response without logged authorized usage (e.g., permission from an access layer), a violation is proven. This can lead to collateral slashing, deterring unauthorized use or distribution, as each distributed model copy contains unique fingerprints. OML 1.0 prioritizes efficiency and offers "next-day security," where compliance is incentivized through the high likelihood of post-violation detection and significant penalties, inspired by optimistic security principles. Further details on OML 1.0’s workflow, specific implementation, and experimental results are elaborated in Appendix B.
Beyond individual model protection with OML primitives like OML 1.0, fostering an incentive-aligned ecosystem requires a broader framework. We conceptually outline a multi-layered, blockchain-based end-to-end protocol designed for this purpose, and a detailed exploration of this end-to-end framework, including its workflow and security analysis, is available in Appendix C.
4 Alternative Views
In this section, we explore alternative perspectives that question the necessity and feasibility of OML, and we articulate our vision by critically engaging with these viewpoints.
4.1 Necessity: Are Existing Approaches Sufficient?
Some may argue OML is unnecessary, as existing paradigms may already form an efficient market equilibrium. OpenAI’s black-box API model (1.2) offers convenience, rapid iteration, and straightforward monetization, appealing to users prioritizing ease and performance. Conversely, HuggingFace’s fully open-weight approach (1.2) champions transparency, reproducibility, and community-driven innovation, allowing extensive local customization. These distinct models currently cater to different, significant segments of the AI landscape, raising questions on the need of a new framework.
4.2 Feasibility: Economic and Technical Concerns
Feasibility concerns also challenge OML. Theoretical OML constructions discussed in Section 3.1, such as those using advanced cryptography or TEEs, can introduce significant cost and latency, potentially hindering democratized AI access. More practical solutions like OML 1.0, while reducing overheads via fingerprinting, offer weaker, post-hoc security guarantees, and may not prevent sophisticated attacks. Furthermore, OML does not fully eliminate risks from malicious model creators embedding hidden backdoors prior to model release to create biased or downgraded responses. Finally, some may argue that policy-based alternative, like regulatory audits and licensing, could more effectively address misuse and ethical concerns without technical overheads inherent to OML.
4.3 Our Vision
Acknowledging the distinct strengths of current paradigms: closed-source APIs (e.g., OpenAI), fully open-weight platforms (e.g., Hugging Face), and policy-driven frameworks, our vision for OML engages these as follows:
  • While closed systems and policy approaches have roles, over-reliance on dominant corporations or centralized mandates risks fragile, distorted AI futures. OML counters by promoting an idealized ecosystem with open accessibility and robust monetizability, fostering collective management and sustainable rewards to prevent singular dominance.
  • While the open-weight paradigm fuels collaborative innovation, it often struggles with sustainable funding for foundational model creation. OML aims for economic self-sufficiency, empowering foundational model builders, not just application developers, to continuously produce and refine high-quality models.
We recognize significant open questions regarding technical limitations and creator-related vulnerabilities, such as backdoor risks (pertinent to many open-weight models), the substantial challenges in balancing cost, performance, and security for OML constructions, and the hardness of finding and embedding robust and undetectable fingerprints into AI models. Precisely due to these complexities, this paper serves as a Call for Extensive Research into more robust, efficient, user-friendly, and secure OML designs. By outlining OML’s conceptual framework, we aim to stimulate broader inquiry into its theory, feasibility, and viability. Our aspiration is for OML to cornerstone a fair, resilient, self-sustaining AI ecosystem, paving the way for equitable benefits.
5 Conclusion
In this position paper, we introduce the Open-access, Monetizable, and Loyal (OML) paradigm as a framework crucial for the fair distribution, sustainable deployment, and equitable governance of AI models. The paper underscores the necessity for OML (Section 1), formalizes its core properties and definitions (Section 2), and explores theoretical constructions alongside practical pathways, including OML 1.0 and a supporting ecosystem (Sections 3). By proposing OML, we aim to catalyze extensive research and collaborative initiatives towards developing next-generation AI systems that are open, resilient, and foster a truly equitable and beneficial AI future for all.
Appendix A Canonical OML Constructions
A.1 Obfuscation
@@
By combining all these techniques, we can come up with a clear construction for OML (Figure 4).
Figure 4: OML formatting process of AI models via obfuscation.
OML formatting. Recall that a naive OML file can be constructed simply by prepending the permission string verification function ${\rm Verify}_{\rm pk}$ to the plain-text model $M$, with the model only returning the correct result if the verification passes. This implies that an attacker can easily find and remove the verification function in the code, recovering the use of the model without the need for permission. To safeguard this OML construction, software obfuscation techniques can be applied such that the two components (${\rm Verify}_{\rm pk}$ and $M$) are intermingled with one another, represented as non-comprehensible code with complicated control flow. As a result, it is difficult to pinpoint the exact location of ${\rm Verify}_{\rm pk}$ in the obfuscated OML file, making it hard for an attacker to remove verification and recover the original model $M$.
@@
A.2 Fingerprinting
Optimistic OML prioritizes efficiency while ensuring a weaker notion of next-day security, i.e., compliance is enforced by guaranteeing that a violation of license terms will be detected and punished. Inspired by optimistic security [81], optimistic OML relies on compliance with the license, and compensating transactions are used to ensure that the model owners’ rights are protected, in case of a violation. Crucial in this process are techniques for authenticating the ownership of a model. For example, Llama models [104] are released under a unique license that a licensee with more than 700 million monthly active users is “not authorized to exercise any of the rights under this Agreement unless or until Meta otherwise expressly grants you such rights". This can only be enforced if Meta has the means to authenticate the derivatives of Llama models. We propose planting a backdoor on the model such that it memorizes carefully chosen fingerprint pairs of the form (key, target response). If successful, such fingerprints can be checked after deployment to claim ownership. An optimistic OML technique should satisfy the following criteria: